# Local Dissipativity Analysis of Nonlinear Systems

### v1

@@ Abstract
[…]
@@
robust control, nonlinear systems, dissipativity
@@ 1 Introduction
For linear systems, we can establish IO properties through frequency-domain analysis [brogliato2007dissipative], solving LMIs that arise from Hamilton Jacobi Inequalitys [Gupta1996], or data-driven analysis [koch2021provably]. While IO theory takes an operator theoretic approach to dynamical systems analysis [zames1966input, vidyasagar1981input, moylan1978stability], the alternative state-space perspective on QSR dissipativity from [willems1972dissipative, hill1976stability] provides an avenue for analysis by linking a systems internal and supplied energy through a so-called storage function. For linear systems, this relationship led to the celebrated Kalman-Yakubovich-Popov (KYP) Lemma [brogliato2007dissipative], which provides necessary and sufficient conditions for dissipativity. The KYP Lemma equates dissipativity with the existence of a positive, semi-definite, quadratic storage function, which can be found by solving LMIs via optimization. Analogous to linear systems, dissipativity of a nonlinear, control affine system is equivalent to the existence of a positive semi-definite storage function [hill1976stability, hill1980connections]. However, because the storage function lacks a set functional form, there is no systematic way to find it, so model-dependent ad-hoc ingenuity is needed. As a result, much effort has been focused on developing model agnostic analysis methods for nonlinear systems.
IO analysis of nonlinear systems requires synthesis of the system’s storage function that adheres to the associated HJI. The problem of IO analysis of nonlinear systems can be approached from an optimal control perspective – using dynamic programming (DP) schemes to approximate a system’s storage function [james1995numerical, dower2008dynamic]. However, [dower2008dynamic] is limited to determining gain over a finite horizon, while [james1995numerical] uses iterative methods that require a good initialization and do not always have convergence guarantees. Another approach is to transform the nonlinear system into an easier form for analysis [sutavani2023small, zahedzadeh2009upper]. For example, in [sutavani2023small], nonlinear systems are approximated using Koopman operators so that analysis tools for linear systems can be used [sutavani2023small]. These methods are only valid in a limited region of the state space where the transformation is valid, and often require ad hoc ingenuity to determine the transformation.
@@
Recent work formulated a nonconvex optimization problem that bounds the gain a of a nonlinear system by searching for a CPA storage function that satisfies the HJI [lavaei2022simultaneous] Constraints were imposed on the storage function at each vertex of the triangulation to ensure certain properties, leveraging a similar process as in previous work on Lyapunov function synthesis [giesl2012construction, giesl2014revised]. Specifically, the optimization problem in [lavaei2022simultaneous] depends on a specific error term that, when imposed with the HJI on each vertex of the triangulation, ensures the HJI holds for the entire region. However, the HJI and its error term are polynomial in design variables – leading to a conservative, nonconvex optimization problem that requires methods like IO to solve. Moreover, specific requirements on the error bound restricted the optimization problem to a limited class of nonlinear systems – notably precluding linear and control affine terms.
@@
This paper determines the local QSR dissipativity characterization of nonlinear, control affine systems via synthesis of a CPA storage function that adheres to the HJI over a region of the state space. This analysis is accomplished via a convex optimization problem that simultaneously searches for the QSR parameters and CPA storage function that satisfies that the HJI – ensuring that the inequality holds for the entirety of the region. In essence, our approach triangulates a region of the state space, and then upper bounds the nonlinear dynamics of the system on each simplex with a linear system. We are then able to synthesize a storage function across the linearized systems.
In Section 3, we develop a novel LMI error bound to impose LMI constraints on a CPA function on a triangulation. We also develop a method to enforce inequalities on a closed norm ball about the origin, which allows us to consider quadratic storage functions. The development of these error bounds builds on work in [strong2024improved]. In Section 4, we develop conditions for a CPA storage function to adhere to the HJI. Because of the LMI error bounds, we are able to reformulate the HJI as an LMI– convexifying the previous optimization problem. Optimization problems are formulated to search for general dissipativity characterizations of nonlinear systems. In Section 5 theoretical results are developed to confirm that if a nonlinear control affine system is strictly locally dissipative, then this optimization problem can be used to verify dissipativity characteristics for either CPA or quadratic storage functions. Finally, numerical results are given in Section 6.
@@
For the function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n},$ let $f^{(k)}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ denote the $k^{th}$ element of $f,$ and let $\mathbf{H}_{f^{(k)}}$ denote the Hessian for this element, where $k\in\mathbb{Z}_{1}^{n}.$ Further, let $f{\in}\mathcal{C}^{r}(\mathcal{X})$ denote a real valued function is $r$-times differentiable over the domain $\mathcal{X}$. The right-hand (left-hand) upper Dini-derivatives for a function, $f$, is defined $D^{+}f(\mathbf{y})\vcentcolon=\lim\sup_{k\rightarrow 0^{+}}\frac{f(\mathbf{x}{+}kg(\mathbf{x})){-}f(\mathbf{x})}{k}$
$\left(\lim\sup_{k\rightarrow 0^{-}}\frac{f(\mathbf{x}{+}kg(\mathbf{x})){-}f(\mathbf{x})}{k}\right),$ where $k{\in}\mathbb{R}$ and $\dot{\mathbf{x}}{=}g(\mathbf{x})$ [giesl2014revised].
@@ 2 Problem Statement and Background
Here, $\bar{G}(\mathbf{x})=\mathbf{B}+G(\mathbf{x})$ and $\bar{J}(\mathbf{x})=\mathbf{D}+J(\mathbf{x}).$ Let $f(0)=0,$ $G(0)=0,$ $h(0)=0,$ and $J(0)=0$.
We will establish local dissipativity by developing constraints on CPA storage functions implying QSR dissipativity. Local dissipativity will be sought for a family of mappings, $\mathcal{G}_{\Omega}\vcentcolon=\{\mathcal{G}(x_{0})\mid x_{0}\in\Omega\},$ defined over a robustly positive invariant subset of the state space, $\Omega\subset\mathbb{R}^{n}$, for inputs in a non-empty set $\mathcal{U}_{\Omega}\subseteq\mathcal{L}_{2}^{m}$, where $\mathcal{U}_{\Omega}\vcentcolon=\{\mathcal{U}_{\Omega}(x_{0})\mid x_{0}\in\Omega\}.$ Here, if $x_{0}\in\Omega$ and $\mathbf{u}\in\mathcal{U}_{\Omega}(x_{0})$, then $x(t)\in\Omega$ for all $t\geq 0$.
@@ 2.1 Local Dissipativity
[willems1972dissipative, Definition 2]: A dynamical system (1) with a locally integrable supply rate $w(t)=w(u,y)$ is dissipative, if there exists a non-negative function, $V:\mathbb{R}^{n}\rightarrow\mathbb{R}_{+}$, called the storage function, such that for all $(t_{0},t_{1})\in\mathbb{R}_{+}$, $x\in\mathbb{R}^{n}$, and $\mathbf{u}\in\mathcal{L}_{2}^{m}$
@@
    The set $\mathcal{U}_{\Omega}(x_{0})$ is non-empty, full dimensional, and contains a neighborhood of the origin.
@@
The family of mappings $\mathcal{G}_{\Omega}$ is locally dissipative if all of its elements are.
The quadratic supply rate in (3) admits sufficient conditions for QSR dissipativity of smooth, nonlinear, control affine systems [hill1976stability, Theorem 1], [van19922, Theorem 2].However, these were originally formulated for global dissipativity. Local dissipativity can be established instead by applying the proof strategy from [van19922, Theorem 2] while restricting the initial state to a set, $x_{0}\in\Omega$, that is robustly positive invariant to inputs in $u\in\mathcal{U}_{\Omega}\neq\emptyset$ and using a more general storage function. The resulting conditions implying local dissipativity of (1) in $\Omega$ when $\mathcal{U}_{\Omega}$ are summarized next, providing an LMI that mirrors the KYP Lemma.
@@
[hill1976stability, Theorem 1], [van19922, Theorem 2] Consider (1) on $\Omega\subset\mathbb{R}^{n}$, where $\mathcal{U}_{\Omega}(x_{0})$ is non-empty for all $x_{0}\in\Omega$. If there exists a non-negative, smooth function, $V:\Omega\rightarrow\mathbb{R}_{+},$ that satisfies
@@
Let $x_{0}\in\Omega$ and $\mathbf{u}\in\mathcal{U}_{\Omega},$ then [hill1976stability, Equation 6] holds where $V(x)=\phi(x).$ Thus, $\mathcal{G}_{\Omega}$ adheres to [hill1976stability, Theorem 1] for $x_{0}\in\Omega$ and $\mathbf{u}\in\mathcal{U}.$
@@ 2.2 Continuous Piecewise Affine Functions
[giesl2014revised, Remark 9] Consider the triangulation $\mathcal{T}=\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}},$ where $\sigma_{i}=\text{co}(\{x_{i,j}\}_{j=0}^{n})$, and a set $\mathbf{W}=\{W_{x}\}_{\mathbf{x}\in\mathbb{E_{\mathcal{T}}}}\subset\mathbb{R},$ where $W(x)=W_{x},\forall x\in\mathbb{E}_{\mathcal{T}}.$ For simplex $\sigma_{i}$, let $\mathbf{X}_{i}\in\mathbb{R}^{n\times n}$ be a matrix that has $x_{i,j}-x_{i,0}$ as its $j$-th row and $\bar{W}_{i}\in\mathbb{R}^{n}$ be a vector that has $W_{x_{i,j}}-W_{x_{i,0}},$ as its $j$-th element. The function $W(x)=x^{\top}\mathbf{X}_{i}^{-1}\bar{W}_{i},$ is the unique CPA interpolation of $\mathbf{W}$ on $\mathcal{T}$ for $x\in\sigma_{i}$.
Like in [giesl2012construction], Taylor’s theorem (Theorem 8) is used to bound the change between a function $\zeta\in\mathcal{C}^{2}$ when evaluated at an arbitrary point versus the vertices on a simplex.
@@
Consider $\zeta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, where $\zeta\in\mathcal{C}^{2}.$ For any $x,y\in\mathbb{R}^{n},$
@@
Although storage functions must satisfy more complicated constraints than Lyapunov functions, the error bounds used to establish CPA Lyapunov functions in [giesl2012construction, giesl2014revised] be useful here.
@@
In Section A, Lemma 17 derives a similar, but tighter, error bound than the one presented above. The proof also uses Taylor’s theorem to bound the change in a function across a simplex. But while Lemma 9 performs a Taylor expansion about a vertex point of the simplex, $x_{0}$, Lemma 17 (Section A) performs the expansion about an arbitrary point within the simplex. Importantly, the behavior of these error bounds differs at the origin. When $x_{j}=0$, $c_{j}=0$ in Lemma 9, while $c_{j}\neq 0$ in Lemma 17. Thus, we will use the tighter bound from Lemma 17 throughout, except when simplices contain the origin. Then, we will revert to using Lemma 9.
@@ 3 Error bounds for function synthesis
[…]
Additionally, in Section 3.2, an LMI to enforce a scalar inequality constraint on a closed ball about the origin is developed. This result is later used for synthesis of combined CPA and quadratic storage functions in dissipativity analysis.
@@ 3.1 LMI Error Bounds
A positive definite error bound matrix is developed for an LMI constraint applied to an $n$-simplex. Enforcing the LMI constraint plus its error bound on the vertex points of an $n$-simplex ($\sigma=\text{co}\{x_{j}\}_{j=0}^{n}$) implies that the LMI holds for all points within that simplex ($x\in\sigma$). While [giesl2012construction] established Lemma 9 as a general error bound for $\mathcal{C}^{2}$ vector-valued functions, it does not translate automatically to an LMI that contains $\mathcal{C}^{2}$ vector-valued functions because the structure of an LMI affects its definiteness. This factor is considered in the following theorem.
@@
Theorem 10 bounds the difference between $\mathbf{M}(x)$ at any convex combination of vertex points, $\mathbf{M}(\sum_{i=0}^{n}\lambda_{i}\mathbf{x}_{i})$, and the convex combination of the LMI evaluated at each vertex point, $\sum_{{j}=0}^{n}\lambda_{j}\mathbf{M}(x_{j})$. The proof parallels [giesl2012construction, Proposition 2.2] by developing remainder terms using Taylor’s theorem, but exploits the structure of $\mathbf{M}(x)$ to establish an LMI error bound. Note the negative identity matrix in (10) is essential to enforce negative semi-definiteness.
@@
In Theorem 10, functions contained within then LMI are linearized about an arbitrary point within a simplex ($x\in\sigma$). However, a similar process as to Lemma 9 can used to linearize the functions about a fixed vertex, $x_{0}.$ The resulting corollary of this process is stated in Section B.

### v2

@@ Abstract
Dissipativity is an input-output (IO) characterization of nonlinear systems that enables compositional robust control through Vidyasagar’s Network Dissipativity Theorem (VNDT). However, determining the dissipativity of a system is an involved and, often, model-specific process. We present a general method to determine the local dissipativity properties of smooth, nonlinear, control affine systems. We simultaneously search for the optimal IO characterization of a system and synthesize a continuous piecewise affine (CPA) storage function via a convex optimization problem. To do so, we reformulate the dissipation inequality as a matrix inequality (MI) and develop novel linear matrix inequality (LMI) bounds for a triangulation to impose the dissipativity conditions on the CPA storage function Further, we develop a method to synthesize a combined quadratic and CPA storage function to expand the systems the optimization problem is applicable to. Finally, we establish that our method will always find a feasible IO characterization and storage function given that the system is sufficiently strictly locally dissipative and demonstrate the efficacy of our method in determining the conic bounds and gain of various nonlinear systems.
@@
dissipativity, analysis of nonlinear systems, robust control of nonlinear systems
@@ 1 Introduction
For linear systems, we can establish IO properties through frequency-domain analysis [brogliato2007dissipative], solving LMIs that arise from Hamilton Jacobi Inequalities [Gupta1996], or data-driven analysis [koch2021provably]. While IO theory takes an operator theoretic approach to dynamical systems analysis [zames1966input, vidyasagar1981input, moylan1978stability], the alternative state-space perspective on QSR dissipativity from [willems1972dissipative, hill1976stability] provides an avenue for analysis by linking a systems internal and supplied energy through a so-called storage function. For linear systems, this relationship led to the celebrated Kalman-Yakubovich-Popov (KYP) Lemma [brogliato2007dissipative], which provides necessary and sufficient conditions for dissipativity. The KYP Lemma equates dissipativity with the existence of a positive, semi-definite, quadratic storage function, which can be found by solving LMIs via optimization. Analogous to linear systems, dissipativity of a nonlinear, control affine system is equivalent to the existence of a positive semi-definite storage function [hill1976stability, hill1980connections]. However, because the storage function lacks a set functional form, there is no systematic way to find it, so model-dependent ad-hoc ingenuity is needed. As a result, much effort has been focused on developing model agnostic analysis methods for nonlinear systems.
IO analysis of nonlinear systems requires synthesis of the system’s storage function that adheres to the associated HJI. The problem of IO analysis of nonlinear systems can be approached from an optimal control perspective – using dynamic programming (DP) schemes to approximate a system’s storage function [james1995numerical, dower2008dynamic]. However, [dower2008dynamic] is limited to determining gain over a finite horizon, while [james1995numerical] uses iterative methods that require a good initialization and do not always have convergence guarantees. Another approach is to transform the nonlinear system into an easier form for analysis [sutavani2023small, zahedzadeh2009upper]. For example, in [sutavani2023small], nonlinear systems are represented using Koopman operators so that analysis tools for linear systems can be used [sutavani2023small]; however, this method is only applicable in a limited region of the state space where the transformation is valid. On the other hand, transformations in [zahedzadeh2009upper] require specific classes of nonlinear systems.
@@
Preliminary work tackled the special case of identifying gain of nonlinear systems by formulating the search for a CPA storage function as convex optimization problems [lavaei2022simultaneous, strong2024improved]. Constraints were imposed on the storage function at each vertex of the triangulation to ensure certain properties, leveraging a similar process as in previous work on Lyapunov function synthesis [giesl2012construction, giesl2014revised]. We aim to expand upon this approach, admitting nonzero control affine terms where [lavaei2022simultaneous] did not, and examining QSR dissipativity instead of gain.
@@
This paper determines the local QSR dissipativity characterization of nonlinear, control affine systems via synthesis of a CPA storage function that adheres to the dissipation condition over a region of the state space. This analysis is accomplished via a convex optimization problem that simultaneously searches for the QSR parameters and CPA storage function that satisfies that the dissipation condition – ensuring that the inequality holds for the entirety of the region. In essence, our approach triangulates a region of the state space, and then upper bounds the nonlinear dynamics of the system on each simplex with a linear system. We are then able to synthesize a storage function across the linearized systems.
In Section 3, we develop a novel LMI error bound to impose MI constraints on a CPA function on a triangulation. We also develop a method to enforce inequalities on a closed norm ball about the origin, which allows us to consider quadratic storage functions. The development of these error bounds builds on work in [strong2024improved]. In Section 4, we develop conditions for a CPA storage function to adhere to the dissipation condition. Because of the LMI error bounds, we are able to reformulate the dissipation condition as an MI– convexifying the previous optimization problem. Optimization problems are formulated to search for general dissipativity characterizations of nonlinear systems. In Section 5 theoretical results are developed to confirm that if a nonlinear control affine system is sufficiently strictly locally dissipative, then this optimization problem can be used to verify dissipativity characteristics. Finally, numerical results are given in Section 6.
@@
For the function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n},$ let $f^{(k)}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ denote the $k^{th}$ element of $f,$ and let $\mathbf{H}_{f^{(k)}}$ denote the Hessian for this element, where $k\in\mathbb{Z}_{1}^{n}.$ Further, let $f{\in}\mathcal{C}^{r}(\mathcal{X})$ denote a real valued function is $r$-times differentiable over the domain $\mathcal{X}$. The right-hand (left-hand) upper Dini-derivatives for a function, $f$, is defined $D^{+}f(\mathbf{y})\vcentcolon=\lim\sup_{k\rightarrow 0^{+}}\frac{f(x{+}kg(x)){-}f(x)}{k}$
$\left(\lim\sup_{k\rightarrow 0^{-}}\frac{f(x{+}kg(x)){-}f(x)}{k}\right),$ where $k{\in}\mathbb{R}$ and $\dot{x}{=}g(x)$ [giesl2014revised].
@@ 2 Problem Statement and Background
Here, $\bar{G}(x)=\mathbf{B}+G(x)$ and $\bar{J}(x)=\mathbf{D}+J(x).$ Let $f(0)=0,$ $G(0)=0,$ $h(0)=0,$ and $J(0)=0$.
We will establish local dissipativity by developing constraints on CPA storage functions implying QSR dissipativity. Local dissipativity will be sought for a family of mappings, $\mathcal{G}_{\Omega}\vcentcolon=\{\mathcal{G}(x_{0})\mid x_{0}\in\Omega\},$ defined over a robustly positive invariant subset of the state space, $\Omega\subset\mathbb{R}^{n}$, for inputs in a non-empty set $\mathcal{U}_{\Omega}\subseteq\mathcal{L}_{2}^{m}$ defined by
$\mathcal{U}_{\Omega}\vcentcolon$= $\bigcap_{x_{0}\in\Omega}\mathcal{U}_{\Omega}(x_{0})$
$\mathcal{U}_{\Omega}(x_{0})\vcentcolon$= $\left\{\mathbf{u}\in\mathcal{L}_{2e}^{m}\mid x(t)\in\Omega\ \forall t\geq 0\right\}.$
@@ 2.1 Local Dissipativity
[willems1972dissipative, Definition 2]: A dynamical system (1) is dissipative with respect to a locally integrable supply rate $w(t)=w(u,y)$ if there exists a non-negative function, $V:\mathbb{R}^{n}\rightarrow\mathbb{R}_{+}$, called the storage function, such that for all $(t_{0},t_{1})\in\mathbb{R}_{+}$, $x\in\mathbb{R}^{n}$, and $\mathbf{u}\in\mathcal{L}_{2}^{m}$
@@
    The set $\mathcal{U}_{\Omega}(x_{0})$ contains a neighborhood of the origin.
@@
The family of mappings $\mathcal{G}_{\Omega}$ is locally QSR dissipative on $\Omega$ with inputs in $\mathcal{U}_{\Omega}$ if all of its elements are dissipative with a shared supply rate and $\mathcal{U}_{\Omega}$ contains a neighborhood of the origin.
Sufficient conditions for global QSR dissipativity of smooth, nonlinear, control affine systems were formatted in [hill1976stability, Theorem 1], [van19922, Theorem 2]. Local dissipativity can be established instead by applying the proof strategy from [van19922, Theorem 2] while restricting the initial state to a set, $x_{0}\in\Omega$, that is robustly positive invariant to inputs in $u\in\mathcal{U}_{\Omega}\neq\emptyset$ and using a more general storage function. The resulting conditions are summarized next, providing a MI that mirrors the KYP Lemma.
@@
[hill1976stability, Theorem 1], [van19922, Theorem 2] Suppose that for some $\Omega\subset\mathbb{R}^{n}$ and dynamics defined by (1), $\mathcal{U}_{\Omega}$ contains a neighborhood of the origin. If there exists a non-negative, smooth function, $V:\Omega\rightarrow\mathbb{R}_{+},$ that satisfies
@@
First note that for each $x_{0}\in\Omega$, $\mathcal{U}_{\Omega}(x_{0})$ contains a neighborhood of the origin because $\mathcal{U}_{\Omega}\subseteq\mathcal{U}_{\Omega}(x_{0})$. For $\mathbf{u}\in\mathcal{U}_{\Omega}(x_{0}),$ we can follow the steps proving [hill1976stability, Theorem 1] with $V(x)=\phi(x)$ up to [hill1976stability, Equation 6], which verifies (4), and consequently local QSR dissipativity.
@@ 2.2 Continuous Piecewise Affine Functions
[giesl2014revised, Remark 9] Consider the triangulation $\mathcal{T}=\{\sigma_{i}\}_{i=1}^{m_{\mathcal{T}}},$ where $\sigma_{i}=\text{co}(\{x_{i,j}\}_{j=0}^{n})$, and a set $\mathbf{W}=\{W_{x}\}_{x\in\mathbb{E_{\mathcal{T}}}}\subset\mathbb{R},$ where $W(x)=W_{x},\forall x\in\mathbb{E}_{\mathcal{T}}.$ For simplex $\sigma_{i}$, let $\mathbf{X}_{i}\in\mathbb{R}^{n\times n}$ be a matrix that has $x_{i,j}-x_{i,0}$ as its $j$-th row and $\bar{W}_{i}\in\mathbb{R}^{n}$ be a vector that has $W_{x_{i,j}}-W_{x_{i,0}},$ as its $j$-th element. The function $W(x)=x^{\top}\mathbf{X}_{i}^{-1}\bar{W}_{i},$ is the unique CPA interpolation of $\mathbf{W}$ on $\mathcal{T}$ for $x\in\sigma_{i}$.
Like in [giesl2012construction], Taylor’s Theorem is used to bound the change between a function $\zeta\in\mathcal{C}^{2}$ when evaluated at an arbitrary point versus the vertices on a simplex.
@@
Consider
$\zeta:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, where $\zeta\in\mathcal{C}^{2}.$ For any $x,y\in\mathbb{R}^{n},$
@@
Although storage functions must satisfy more complicated constraints than Lyapunov functions, the error bounds used to establish CPA Lyapunov functions in [giesl2012construction, giesl2014revised] will be useful here.
@@
In Appendix A, Lemma 18 establishes a similar, but tighter, error bound than the one presented above. The proof also uses Taylor’s Theorem to bound the change in a function across a simplex. But while Lemma 9 performs a Taylor expansion about a vertex point of the simplex, $x_{0}$, Lemma 18 (Appendix A) performs the expansion about an arbitrary point within the simplex. Importantly, the behavior of these error bounds differs at the origin. When $x_{j}=0$, $c_{j}=0$ in Lemma 9, while $c_{j}\neq 0$ in Lemma 18. Thus, we will use the tighter bound from Lemma 18 throughout, except when simplices contain the origin. Then, we will revert to using Lemma 9.
@@ 3 Error bounds for function synthesis
Lemma 9 is an invaluable tool for ensuring that CPA function adheres to a scalar inequality condition. In this section, we expand this tool to more complex MI conditions, which often codify complex control analysis and synthesis requirements. We develop novel error bounds for an MI on an n-simplex and later leverage them for dissipativity analysis.
In Section 3.2, an additional LMI is developed to enforce a scalar inequality constraint over a closed ball about the origin. This is later used to combine CPA and quadratic storage functions for dissipativity analysis.
@@ 3.1 LMI Error Bounds
A positive definite error bound matrix is developed for a MI constraint applied to an $n$-simplex. Enforcing the MI constraint plus its error bound on the vertex points of an $n$-simplex ($\sigma=\text{co}\{x_{j}\}_{j=0}^{n}$) implies that the MI holds for all points within that simplex ($x\in\sigma$). While [giesl2012construction] established Lemma 9 as a general error bound for $\mathcal{C}^{2}$ vector-valued functions, it does not translate automatically to a MI that contains $\mathcal{C}^{2}$ vector-valued functions because the structure of a MI affects its definiteness. This factor is considered in the following theorem.
@@
Theorem 10 bounds the difference between $\mathbf{M}(x)$ at any convex combination of vertex points, $\mathbf{M}(\sum_{i=0}^{n}\lambda_{i}\mathbf{x}_{i})$, and the convex combination of it evaluated at each vertex point, $\sum_{{j}=0}^{n}\lambda_{j}\mathbf{M}(x_{j})$. The proof parallels [giesl2012construction, Proposition 2.2] by developing remainder terms using Taylor’s Theorem, but exploits the structure of $\mathbf{M}(x)$ to establish an LMI error bound. Note the negative identity matrix in (10) is essential to enforce negative semi-definiteness.
@@
In Theorem 10, functions contained within the MI are linearized about an arbitrary point within a simplex ($x\in\sigma$). However, a similar process to Lemma 9 can used to linearize the functions about a fixed vertex, $x_{0},$ resulting in the corollary stated in Appendix B.